OpenStudy (anonymous):

75 cents/h= dollars/day

sammixboo (sammixboo):

\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm W}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm L}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm C}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm M}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm E}}}\) \(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ black}{\large {\bbox[5pt, white ,border:2px solid white ]{ \rm O}}}\) \(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm O}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm P}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm E}}}\)\(\large\color{ white}{\large {\bbox[5pt,#00c5ff ,border:2px solid white ]{ \rm N}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm S}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm T}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm U}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm D}}}\)\(\large\color{ white}{\large {\bbox[5pt,#92dba2 ,border:2px solid white ]{ \rm Y}}}\) I see what you have posted is \(\Large\rm \frac{75~cents}{h}=\frac{dollars}{day}\), but what is your full problem? I can try to give you the best help I can :)

OpenStudy (jdoe0001):

\(\bf \cfrac{75\cancel{ c }}{\cancel{ hr }}\cdot \cfrac{24\cancel{ hr }}{day}\cdot \cfrac{dollar}{100\cancel{ c }}\) keep in mind that, there are 24hrs/day and 100cents/dollar